OpenStudy (anonymous):

In 2007, the soccer club held a fundraiser and made a profit of $459. This was 85%of thier revenue. The revenue in 2007 was 90% of the revenue from the club's 2008 fundraiser.What was the revenue in 2008? please solve and show work.

OpenStudy (anonymous):

0.85R = 459, so R = 459/0.85 R = 540

OpenStudy (anonymous):

For 2008, we'll call the revenue S. So, 540 = 0.9S S = 540/0.9 S = 600

OpenStudy (anonymous):

ad then waht do u do or thats it?

OpenStudy (anonymous):

Depends on how much work you have to show. But the question asks what was the revenue in 2008?, so that would be the value of S. The revenue in 2008 was $600.

OpenStudy (anonymous):

ahhh is see